OpenStudy (anonymous):

find value of x......... 4^x-3^[x-(1/2)] = 3^[x+(1/2)]-2^(2x-1)

OpenStudy (anonymous):

4^x-3^[x-(1/2)] = 3^[x+(1/2)]-2^(2x-1) 4^x -3^x 3^(-1/2)=3^x * 3^1/2 -2^2x * 2^(-1) 4^x -4^x * 2^(-1)v=3^x * 3^(1/2) + 3^x * 3^(-1/2)

OpenStudy (anonymous):

Does it help?